# 3D-BBS: Global Localization for 3D Point Cloud Scan Matching Using Branch-and-Bound Algorithm

### v4

@@ II-A Global Localization with Point Cloud Matching
Frame-based methods extract geometrical features to represent a whole point cloud frame and identify revisited places by finding frames with similar descriptors. Scan context [11] creates a database of descriptors based on the height of 3D LiDAR points and effectively detects the loops in a trajectory. OverlapNet [12] utilizes a deep neural network exploiting different cues generated from LiDAR data, and estimates the overlap rates and yaw angle between two point clouds. OverlapTransformer [13] uses only LiDAR data, and achieves loop closure candidates detection faster than OverlapNet. BoW3D [14] built a dictionary of bag-of-wards based on Link3D [15] features representing LiDAR scan keypoints and achieved real-time performance in 6DoF loop close. While frame-based methods require keyframes of LiDAR scans that cover the entire map, 3D-BBS directly uses 3D maps that have already been constructed.
@@ III-A Problem Setting
$\bm{x}^{*}=\mathop{\rm argmax}\limits_{\begin{subarray}{c}x\in\{ri_{x}|i_{x}\in\mathbb{Z},W_{\text{min}}^{x}\leq ri_{x}\leq W_{\text{max}}^{x}\}\\ y\in\{ri_{y}|i_{y}\in\mathbb{Z},W_{\text{min}}^{y}\leq ri_{y}\leq W_{\text{max}}^{y}\}\\ \theta\in\{ri_{\theta}|i_{\theta}\in\mathbb{Z},W_{\text{min}}^{\theta}\leq ri_{\theta}\leq W_{\text{max}}^{\theta}\}\end{subarray}}\quad\sum_{k=1}^{K}M(\bm{T}_{\bm{x}}\bm{s}_{k}),$ (4)
@@
Branch-and-bound scan matching [4]: To solve the problem of 2D global localization using the BnB algorithm, the pose space within the search range is represented as a tree structure. Let $l\in\mathbb{Z},0\leq l\leq l_{\text{max}}$ be the hierarchy level in the tree, where $l_{\text{max}}$ is the maximum level. The tree is branched from large to small step sizes. The step size at each level is determined based on the cell size $r$ of the occupancy grid map. 1) The translational step size is $r_{l}=2^{l}r$. 2) The rotational step size $\delta(r)$ is common for each level and selected such that a point in the maximum scan range $d_{\text{max}}$ does not move more than $r$ as follows:
@@ III-D Branching Algorithm
$j_{\beta}\in\{0,...,a_{\beta}-1\},j_{\gamma}\in\{0,...,a_{\gamma}-1\},$ (19)
@@ IV Experiments
To evaluate the estimation result, we used two criteria: 1) the translation error is smaller than 2.0 m, and 2) the rotation error was smaller than 0.05 rad. If a result satisfies both the translation and rotation criteria, we consider it sufficiently accurate to be used as an initial estimate for local fine registration [22] to obtain a fine estimation result.

### v5

@@ II-A Global Localization with Point Cloud Matching
Frame-based methods extract geometrical features to represent a whole point cloud frame and identify revisited places by finding frames with similar descriptors. Scan context [11] creates a database of descriptors based on the height of 3D LiDAR points and effectively detects the loops in a trajectory. OverlapNet [12] utilizes a deep neural network exploiting different cues generated from LiDAR data, and estimates the overlap rates and yaw angle between two point clouds. OverlapTransformer [13] uses only LiDAR data, and achieves loop closure candidates detection faster than OverlapNet. BoW3D [14] built a dictionary of bag-of-wards based on Link3D [15] features representing LiDAR scan keypoints and achieved real-time performance in 6DoF loop closure. While frame-based methods require keyframes of LiDAR scans that cover the entire map, 3D-BBS directly uses 3D maps that have already been constructed.
@@ III-A Problem Setting
$\bm{x}^{*}=\mathop{\rm argmax}\limits_{\begin{subarray}{c}x\in\{ri_{x}|i_{x}\in\mathbb{Z},W_{\text{min}}^{x}\leq ri_{x}\leq W_{\text{max}}^{x}\}\\ y\in\{ri_{y}|i_{y}\in\mathbb{Z},W_{\text{min}}^{y}\leq ri_{y}\leq W_{\text{max}}^{y}\}\\ \theta\in\{\delta(r)i_{\theta}|i_{\theta}\in\mathbb{Z},W_{\text{min}}^{\theta}\leq\delta(r)i_{\theta}\leq W_{\text{max}}^{\theta}\}\end{subarray}}\quad\sum_{k=1}^{K}M(\bm{T}_{\bm{x}}\bm{s}_{k}),$ (1)
@@
Branch-and-bound scan matching [4]: To solve the problem of 2D global localization using the BnB algorithm, the pose space within the search range is represented as a tree structure. Let $l\in\mathbb{Z},0\leq l\leq l_{\text{max}}$ be the hierarchy level in the tree, where $l_{\text{max}}$ is the maximum level. The tree is branched from large to small step sizes. The step size at each level is determined based on the cell size $r$ of the occupancy grid map. 1) The step size is $r_{l}=2^{l}r$. 2) The rotational step size $\delta(r)$ is common for each level and selected such that a point in the maximum scan range $d_{\text{max}}$ does not move more than $r$ as follows:
@@ III-D Branching Algorithm
$j_{\beta}\in\{0,...,a_{\beta}-1\},j_{\gamma}\in\{0,...,a_{\gamma}-1\}\},$ (14)
@@ IV Experiments
To evaluate the estimation result, we used two criteria: 1) the translation error was smaller than 2.0 m, and 2) the rotation error was smaller than 0.05 rad. If a result satisfies both the translation and rotation criteria, we consider it sufficiently accurate to be used as an initial estimate for local fine registration [22] to obtain a fine estimation result.